# Sentiment and Emotion-aware Multi-criteria Fuzzy Group Decision Making System

### v1

@@ 1 Introduction
Discussions are one of the main tools people use to make decisions. When a decision process involves a group of people and requires multiple points of view to consider, it is commonly conducted in groups of people. GDM [2] systems are tools and processes that allow a group of people to work together to reach a consensus acceptable to all group members. GDM systems can be useful where multiple perspectives need to be considered to arrive at a decision that is acceptable to all members of the group [3], [4].
The current study presents a novel decision-making framework that integrates a voting system, sentiment analysis from debates, fuzzy inference systems, and consensus evaluation to perform small decision-making—determining the most suitable hotel from a set of alternatives. In our approach, experts have the flexibility to engage in GDM processes using natural language. Using sentiment analysis methods, we determine the degree to which an expert accepts a particular alternative and obtains a preference value. The methodology aims to balance quantitative evaluations with qualitative insights to enhance decision accuracy and participant satisfaction.
@@ 2.1 Fuzzy Approach in Decision Making
Group decision-making (GDM) systems use fuzzy logic [5] and fuzzy sets [6] to help groups make decisions [7] when there is uncertainty or ambiguity since its first appearance in the 1970 [8]. These systems allow group members to express their opinions [9] in a way that considers the decision-making process’s fuzziness and imprecision. Fuzzy sets and logic enable the representation of imprecise or ambiguous data, which can help capture the complex preferences and opinions of people in a group context. A number of works focus on applying the fuzzy approach to DM.
@@ 2.2 Sentiment and Emotion Analysis for Decision Making
People usually use natural language to express their opinions during discussions [3]. Natural Language Processing (NLP), allows experts to use everyday language rather than specific formats or numbers. By extracting valuable information from large text data, NLP supports decision-making processes.
@@ 3.1 Preferences, voting system
Figure 1 illustrates the proposed GDM system.
@@
Let a finite set of alternatives, denoted as $X={x_{1},x_{2},…,x_{n}}$. From the most preferred to the least preferred, these alternatives have to be ranked based on the information provided by a finite set of experts, denoted as $E={e_{1},e_{2},...,e_{m}}$. In addition, we have $F={f_{1},f_{2},...,f_{p}}$, and $Z={z_{1},z_{2},...,z_{m*p}}$, which are the features and the weights of the corresponding features, chosen by each of the experts $E={e_{1},e_{2},...,e_{m}}$. In a GDM system, the goal is to rank the alternatives in $X$ by considering the preference values $P_{k}$ provided by each expert $e_{j}$, where $j$ ranges from 1 to $m$. In addition, we have $W={w_{1},w_{2},...,w_{n*p}}$, which are the weights of the corresponding features for each of the alternatives $X={x_{1},x_{2},…,x_{n}}$. Preference assessments: “-1” - against, “0” - does not matter, and “1” – agreement. $Pref^{E_{j}}(X_{i})$ – preference value of $E_{j}$ (expert) about $X_{i}$ (alternative) found using Eq. 1.
@@ 3.2 Sentiment analysis
Using VADER from the NLTK library, Sentiment analysis provides quantitative insights from participant discussions, enhancing our understanding of their emotional responses and preferences alongside quantitative scores. We integrate emotion scores into the sentiment analysis to get the final preference score.
Preference $\text{Preference Score}=\alpha\cdot\text{Joint Sentiment Score}+\beta\cdot\text{Emotion Score}$ (2)
@@ 3.3 Fuzzy Inference System
Figure 2: Input and Output Fuzzy Sets
@@ 4 Experiment and Results
We set up an experiment where four participants engage in a decision-making process for choosing a hotel travel destination. The goal is to assess how preference analysis influences the decision-making process and outcomes, such as overall satisfaction and agreement among the participants. Experts will decide which alternatives are the most appropriate. Therefore, they should provide their preferences and opinions for a particular set of alternatives shown in Table 2. Four participants provided their preferences for each attribute using a scale from -1 to 1, where -1 indicates a negative preference, 0 indicates indifference, and 1 indicates a positive preference (see Table 3). These preferences were used to establish the initial voting system, as shown in Table 4.
@@
  • Closeness to City (minutes): The duration of the car ride to the city centre.
@@
The preference score for each participant regarding each hotel is calculated by comparing the hotel’s features to the group’s average preferences. The steps are as follows. Continuous features such as price, rating, room area, and closeness to the city are normalized by comparing them to their respective mean values across all hotels by using the Formula $\text{Mean}=\frac{\sum_{i=1}^{n}x_{i}}{n}$, where $n$ is the total number of objects, $i$ from 1 to $n$. To standardize the comparison of hotel attributes, we mapped the meal type attribute to numerical coefficients (0.33 for Breakfast and 1.0 for All Inclusive). For features already in binary form values are taken without normalization.
@@
Average
Preference Score
@@
Emotional tones (happy, angry, surprise, sad, and fear) were detected using the Text2Emotion library, which assigns scores to different emotions, as shown in Table 7.
For each hotel, we calculated the average combined sentiment score. The emotion score was derived by considering the positive and negative emotions separately. It was calculated by taking the maximum of the following feelings: negative emotions included ”angry,” ”sad,” and ”fear,” while positive emotions included ”happy” and ”surprise.” The difference between the positive and negative emotion scores was used to create the composite emotion score. This approach ensures that the most intense positive and negative emotions are considered, reflecting the overall emotional tone of the discussion. These Emotion and Sentiment scores were weighted and combined to form the total sentiment score: $TotalsentimentScore=EmotionScore*0.4+SentimentScore*0.6$. This weighting scheme gives more importance to sentiment analysis while still considering emotional intensity.
@@
Next, we used the Fuzzy Inference system to simulate and compute the total preference score based on given input values for voting preference and sentiment preference, total preference values shown in Table 8. The alternatives are ranked based on the preference levels provided by the experts.
After selecting a hotel, participants provided feedback on their agreement and confidence levels regarding the decision. These values were used to compute a feedback score using another fuzzy inference system. The feedback system evaluated the consistency and confidence of the participants’ opinions, as shown in Table 10. Consensus metrics, including the interquartile range (IQR) and mean consensus level, were used to evaluate the overall agreement among participants, as shown in Table 10. The consensus level was classified as high, medium, or low based on the mean consensus score. This classification helped understand the group’s satisfaction with the decision-making process.
@@ 5 Conclusion
We presented a decision-making framework that effectively integrates voting preferences, debate analysis, and fuzzy inference systems to aid in selecting a hotel. The system’s ability to consider quantitative preferences and qualitative sentiment analysis ensures a balanced and well-informed decision. The high consensus level among participants further validates the proposed methodology. This approach can be applied to various GDM scenarios, providing a reliable method for achieving consensus.

### v2

@@ 1 Introduction
Discussions are one of the main tools people use to make decisions. When a decision process involves a group of people and requires multiple points of view to consider, it is commonly conducted in groups of people. GDM [2] systems are tools and processes that allow a group of people to work together to reach a consensus acceptable to all group members. GDM systems can be useful where multiple perspectives need to be considered to arrive at a decision acceptable to all group members [3], [4].
The current study presents a decision-making framework that integrates a voting system, sentiment analysis from debates, fuzzy inference systems (FIS), and consensus evaluation to perform small decision-making — determining the most suitable hotel from a set of alternatives. In our approach, experts have the flexibility to engage in GDM processes using natural language. Using sentiment analysis methods, we determine the degree to which an expert accepts a particular alternative and obtains a preference value. The methodology aims to balance quantitative evaluations with qualitative insights to enhance decision accuracy and participant satisfaction. We use FIS because of its ability to handle uncertainty and vagueness in human opinions effectively, which makes it particularly suitable because it does not rely heavily on large datasets.
The primary contribution of this paper is introducing a novel decision-making framework that integrates sentiment and emotion analysis into a fuzzy logic-based system, enhancing the effectiveness of group decision-making processes in contexts where emotional factors play a critical role.
The remainder of this paper is organized as follows: Section II reviews the relevant literature, focusing on the fuzzy approach in decision-making and sentiment and emotion analysis. Section III describes the methodology, including the preferences and voting system, sentiment analysis, and the fuzzy inference system. Section IV presents the experiment and results. Finally, Section V concludes the paper and discusses potential future research directions.
@@ 2 Related Work
Integrating sentiment and emotion awareness into multi-criteria fuzzy group decision-making systems is an emerging area of research. Recent papers include the SAR-MCMD method, combining multi-criteria decision-making, deep learning, and sentiment analysis [5], sentiment-driven fuzzy cloud multicriteria model for online product ranking [6], dual fine-tuning-based LSGDM using online reviews [7], an intuitionistic fuzzy model for consensus reaching based on public opinions and expert evaluation [8].
@@ 2.1 Fuzzy Approach in Decision Making
Group decision-making (GDM) systems use fuzzy logic [9] and fuzzy sets [10] to help groups make decisions [11] when there is uncertainty or ambiguity since its first appearance in the 1970 [12]. These systems allow group members to express their opinions [13] in a way that considers the decision-making process’s fuzziness and imprecision. Fuzzy sets and logic enable the representation of imprecise or ambiguous data, which can help capture people’s complex preferences and opinions in a group context. A number of works focus on applying the fuzzy approach to DM.
@@ 2.2 Sentiment and Emotion Analysis for Decision Making
People usually use natural language to express their opinions during discussions [3]. Natural Language Processing (NLP) allows experts to use everyday language rather than specific formats or numbers. By extracting valuable information from large text data, NLP supports decision-making processes.
@@ 3.1 Preferences, voting system
Figure 1 illustrates the proposed Sentiment and Emotion-aware Multi-criteria Fuzzy GDM System system.
@@
Let a finite set of alternatives be $X={x_{1},x_{2},…,x_{n}}$. From the most preferred to the least preferred, these alternatives have to be ranked based on the information provided by a finite set of experts, denoted as $E={e_{1},e_{2},...,e_{m}}$. In addition, we have $F={f_{1},f_{2},...,f_{p}}$, and $Z={z_{1},z_{2},...,z_{m*p}}$, which are the features and the weights of the corresponding features, chosen by each of the experts $E={e_{1},e_{2},...,e_{m}}$. Because the situation in our particular setting comprises a group of friends working together to choose a hotel, the experts are considered equal. As a result of the peer-based decision-making process, every participant’s opinion is given equal weight. In a GDM system, the goal is to rank the alternatives in $X$ by considering the preference values $P_{k}$ provided by each expert $e_{j}$, where $j$ ranges from 1 to $m$. In addition, we have $W={w_{1},w_{2},...,w_{n*p}}$, which are the weights of the corresponding features for each of the alternatives $X={x_{1},x_{2},…,x_{n}}$. Preference assessments: “-1” - against, “0” - does not matter, and “1” – agreement. $Pref^{E_{j}}(X_{i})$ – preference value of $E_{j}$ (expert) about $X_{i}$ (alternative) found using Eq. 1.
@@ 3.2 Sentiment analysis
Using VADER [25] from the NLTK library, Sentiment analysis provides quantitative insights from participant discussions, enhancing our understanding of their emotional responses and preferences alongside quantitative scores. We integrate emotion scores into the sentiment analysis to get the final preference score.
$\text{Preference Score}=\alpha\cdot\text{Joint Sentiment Score}+\beta\cdot\text{Emotion Score}$ (2)
@@ 3.3 Fuzzy Inference System
Figure 2: Input and Output Fuzzy Sets of preference scores
@@ 4 Experiment and Results
We set up an experiment in which four participants make a decision about a hotel travel destination. The goal is to assess how preference analysis influences the decision-making process and outcomes, such as overall satisfaction and agreement among the participants. Experts will decide which alternatives are the most appropriate. Therefore, they should provide their preferences and opinions for a particular set of alternatives shown in Table 2. Four participants provided their preferences for each attribute using a scale from -1 to 1, where -1 indicates a negative preference, 0 indicates indifference, and 1 indicates a positive preference (see Table 3). These preferences were used to establish the initial voting system, as shown in Table 4.
@@
  • Closeness to City (minutes): The duration of the car ride to the city center.
@@
The preference score for each participant regarding each hotel is calculated by comparing the hotel’s features to the group’s average preferences. The steps are as follows. Continuous features such as price, rating, room area, and closeness to the city are normalized by comparing them to their respective mean values across all hotels by using the Formula $\text{Mean}=\frac{\sum_{i=1}^{n}x_{i}}{n}$, where $n$ is the total number of objects, $i$ - from 1 to $n$. To standardize the comparison of hotel attributes, we mapped the meal type attribute to numerical coefficients (0.33 for Breakfast and 1.0 for All Inclusive). For features already in binary form, values are taken without normalization.
@@
Average Preference Score
@@
Emotional tones (happy, angry, surprise, sad, and fear) were detected using the Text2Emotion library [26], which assigns scores to different emotions, as shown in Table 7. For each hotel, we calculated the average combined sentiment score. The emotion score was derived by considering the positive and negative emotions separately. It was calculated by taking the maximum of the following feelings: negative emotions included ”angry,” ”sad,” and ”fear,” while positive emotions included ”happy” and ”surprise.” The difference between the positive and negative emotion scores was used to create the composite emotion score. This approach ensures that the most intense positive and negative emotions are considered, reflecting the overall emotional tone of the discussion. These Emotion and Sentiment scores were weighted and combined to form the total sentiment score: TotalsentimentScore=EmotionScore0.4+SentimentScore0.6TotalsentimentScore=EmotionScore*0.4+SentimentScore*0.6. This weighting scheme gives more importance to sentiment analysis while still considering emotional intensity.
These weights were chosen to give slightly more emphasis to sentiment analysis due to its broader contextual relevance. This weighting scheme ensures that while sentiment analysis plays a dominant role, the emotional intensity is still substantially incorporated. In general, they can be adjusted based on the context or derived from user preferences.
@@
Next, we used the Fuzzy inference system to simulate and compute the total preference score (see Fig. 3(a)) based on given input values for voting preference and sentiment preference (see Fig. 5), total preference values shown in Table 8. The alternatives are ranked based on the preference levels provided by the experts.
After selecting a hotel, participants provided feedback on their agreement and confidence levels regarding the decision (see Fig. 3(b)). These values were computed using another fuzzy inference system to compute a feedback score (see Fig. 5). The feedback system evaluated the consistency and confidence of the participants’ opinions, as shown in Table 10. Our feedback system is non-monotonic, but this behavior is justified and appropriate given the interpretation of ”confidence” as the confidence level in the agreement. Our rules capture the nuanced relationship between agreement, confidence in that agreement, and feedback, leading to outputs that make sense given the inputs. Consensus metrics, including the interquartile range (IQR) and mean consensus level, were used to evaluate the overall agreement among participants, as shown in Table 10. The consensus level was classified as high, medium, or low based on the mean consensus score. This classification helped understand the group’s satisfaction with the decision-making process.
(a) Preference score
(b) Feedback score
Figure 3: 3D graphic representations of a set of possible fuzzy system solutions
@@ 5 Conclusion
We presented a novel decision-making framework integrating voting preferences, debate analysis, and fuzzy inference systems. The main contribution of this paper is integrating sentiment and emotion analysis into a fuzzy logic-based system, thereby making traditional group decision-making processes emotion-aware, particularly in situations where emotional factors are crucial. An example of the application of our framework is selecting a hotel. The high consensus level among participants further validates the proposed methodology. This approach can be applied to various GDM scenarios, providing a reliable method for achieving consensus. As for future works, we aim to explore the system’s applicability in more complex contexts and further refine its components to increase its generalizability. While our current work shows that integrating sentiment and emotion analysis into a multi-criteria fuzzy GDM system is feasible, as demonstrated with the hotel selection example, we plan to explore its effectiveness in more varied and complex decision-making scenarios in future research. We also plan to conduct a comprehensive sensitivity analysis and compare our approach with other decision-making frameworks to evaluate its effectiveness and reliability more rigorously.
The proposed system integrates sentiment and emotion analysis, balancing both quantitative and qualitative inputs, making it well-suited for real-world scenarios like the hotel selection case study. However, its complexity due to multiple components can make implementation challenging. The system’s effectiveness may also vary in different decision-making contexts, especially given the simplicity of the current case study. Additionally, the approach depends heavily on the quality of textual data, which could affect outcomes if the data is incomplete or biased.
In the future, we will explore integrating our framework with conventional multi-criteria decision support techniques like TOPSIS, ELECTRE, and PROMETHEE to enable a thorough comparison and assessment of its efficacy in various decision-making scenarios.